OpenStudy (anonymous):

evaluate the definite integral {2 | 1-x2|dx -2 | absolute value of 1-x2

OpenStudy (anonymous):

is it \[\int\limits_{-2}^{2}|1 - x^2 | dx\]

OpenStudy (anonymous):

\[\int\limits_{-2}^{-1}(x^2 - 1)dx + \int\limits_{-1}^{0}(1 - x^2 ) dx + \int\limits_{0}^{1}(1-x^2)dx + \int\limits_{1}^{2}(x^2 -1)dx\]

OpenStudy (anonymous):

u get it?

OpenStudy (anonymous):

plz explain it

OpenStudy (anonymous):

u know the behavior of the absolute value function dont u?

OpenStudy (anonymous):

u break the function according to its behavior in various intervals and then integrate it piecewise

OpenStudy (anonymous):

Find the volume of the solid that results when the region enclosed by curve on the interval is revolved about the y-axis.

OpenStudy (anonymous):

dunno volumes

OpenStudy (anonymous):

i told u how to integrate